OpenStudy (anonymous):

3x+10=1 need answer !!

OpenStudy (anonymous):

-3

OpenStudy (anonymous):

subtract 10 from both sides: 3x=-9 divide both sides by 3: x=-3

OpenStudy (anonymous):

\[3x+10=1 ; 3x=-9 ; x=-9/3 ; x= -3\]

OpenStudy (anonymous):

3x+10=1 3x =1-10 x = -9/3 x = -3